How do you:
Make a the subject (with steps) of :
3(a+4)=ac+5f

Answer by Theo(13342)

you want to solve for a in terms of the other variables is what i think you mean.
start with:
3(a+4) = ac + 5f
distribute the multiplication to get:
3a + 12 = ac + 5f
subtract 12 from both sides of the equation to get:
3a = ac + 5f - 12
subtract ac from both sides of the equation to get:
3a - ac = 5f - 12
factor out the a to get:
a * (3 - c) = 5f - 12
divide both sides of the equation by (3 - c) to get:
a = (5f - 12) / (3 - x)
that's your solution.